### v10

@@ 1 Introduction and overview
Many economists over the centuries have reasoned that net saving, or equivalently net investment [^1] As reported in national accounts; they differ only by statistical discrepancy. Also see Appendix A., should tend to give equal capital growth. Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)). Roy Harrod (1939) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”. Harrod, and many other economists of his time and since, have focused on growth of output rather than of capital, but have modeled growth of output by first assuming the equivalence of net saving and capital growth, within the warranted rate, and then looking for effects of that capital growth on later output growth.
@@
Those test results from data for net saving/capital ratio are easiest to reconcile in an interpretation that net saving has no effect on capital growth. Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
The data are most easily explained by the present value principle, by production efficiencies enabled through innovation, and by investment in the new technology from depreciation of the technology replaced. Value is created in the mind of the market at the moment when prospective cash flows are discounted. It is created only if the market sees a path, step by step, from the start, to practical realization of those prospective cash flows. Then capital growth arrives when the market first evaluates prospective cash flows, and is realized eventually in physical outcomes insofar as the market has predicted correctly. That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment. [^2] Present value is time-discounted, and affords innovators only enough, at first, for first steps in implementation. As implementation proceeds, present value rises and affords more.
Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up. It is that temporary market advantage to the innovator which explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the explanation in terms of market valuation. This idea will be called “free growth theory” for easy reference.
It predicts only at the largest scales, and only for the private sector. Individuals and groups and even small economies can grow or decline through net saving or dissaving. That possibility is foreclosed only at the scale of all capital and all economies together. The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
@@
Inferences for economic teaching include the obvious ones for growth theory and for net saving in general. They include others as well. One of the central doctrines of the marginalist revolution has held that market realization converges to producer cost, when that cost includes imputed interest on assets owned. Net saving gives producer cost, and falls short of market realization in the presence of technological growth from new ideas. Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but not the aggregate value. In general, economics might consider relying less on book value and net investment, and more on market value and on the power of ideas.
@@ 2 Net saving, capital growth and capital acceleration
$\Delta g(K)=\Delta s^{*}\ ,\quad\text{uta.}$. (3)
$\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative. Division of Eq. (3) by capital acceleration, with rearrangement, gives
@@
Test results for the predictions $\frac{s^{*}}{g(K)}=\theta=1$, under thrift assumptions, appear in Fig. 1 and in Tables 1, 2 and 3. Screens to which they refer are explained in Section 14 below. Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts. They show $\frac{s^{*}}{g(K)}=0.481$, with regression of $s^{*}$ on $g(K)$ found at 0.0771, and $\frac{\Delta s^{*}}{\Delta g(K)}=0.064$, with regression of $\Delta s^{*}$ on $\Delta g(K)$ at 0.0559.
@@
  Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗ (0.0144) Observations 1,826 R2 0.84511 Within R2 0.10669 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$ Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗ (0.0057) Observations 1,574 R2 0.37071 Within R2 0.27233 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$
@@
Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$.
@@ 3 Interpretation of test results
Any explanation must account for the chance that replacement saving will sometimes not be enough to offset depreciation. This shortfall is likeliest in prolonged downturns, where loss of income motivates households to forego net saving, and then to invade capital by putting off maintenance expense and replacement investment in order to protect consumption. Effects will include a positive effect on regression of $s^{*}$ on $g(K)$, and an upward effect on $\frac{\Delta s^{*}}{\Delta g(K)}$, as $s^{*}$ and $g(K)$ decline together in response to downward pressure on income. This interpretation could fit the appearance in Fig. 2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises.
In general, it might be expected that households stabilize consumption by building up reserves through net saving when earnings are higher, and deplete them when earnings are lower, as described in the life cycle theory of saving (Modigliani (1954)) or the permanent income hypothesis (Friedman (1957)). Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\theta$ as does thrift theory, but from an opposite causality.
Although these results cannot exclude the possibility that some net saving has effect, data are easiest to reconcile with free growth theory, where all capital growth is explained by productivity gain. Pending further study, we will proceed on this interpretation.
@@ 5 Solow’s puzzle
Some growth is capital widening, where structures and implements increase in number but do not change in design. Capital widening, however, is practical only so far before glut and diminishing returns set in. Further growth from that point must come from capital deepening, meaning improvements in the design of capital. Solow (1956) noted a kind of middle ground between capital widening and capital deepening in the disembodied growth mentioned earlier; ships carrying coals to Newcastle can raise prospective cash flows, and hence present value, by reversing the business plan. But Solow, who came to conclusions similar to ours from different evidence, puzzled as to how capital growth without net saving could be possible for capital deepening through “embodied” growth, where products of new design are made from plant of new design.
Solow’s Nobel Prize acceptance speech in mentions the embodiment problem: [^4] Solow (1988). The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
@@ 6 Mechanics of free growth
The present value argument just given offers an accounting explanation of embodied growth in terms of sources of funds. The step-by-step mechanical explanation, we suggest, in terms of things rather than dollars, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.555i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
Our findings support those of Piketty and Zucman (2014) and Kurz (2023) as to the market power of innovators to explain capital growth beyond net saving. Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
@@ 11 Net output
The concept of net output is found in our discussions, but not in our equations. Net output means value added. Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital, and that both of these components are unmeasurable by known means. Since we work here from measurable data for net saving and market value capital, we did our best to reason from those two quantities alone.
@@ Appendix B.  Net output with human capital
Human capital is impractical to measure, as it leaves little market record other than for its rental income in pay and investment cost in schooling. Thus national accounts leave it implicit, and allow us to infer what we can from data for pay and schooling. Those accounts are founded on the principle, sound when terms are appropriately redefined, that net output, or value added, is expressed in the sum of capital growth and net outflow from the value-added chain. In national accounts, then, where physical capital is the whole of capital while net outflow of the chain is the whole of consumption, the reasoning is
$Y=\Delta K+C\ ,\quad\text{neglecting human capital.}$. (B.1)
It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq. (B.1). Let human capital $H$, in that new model, stand as the last link in the value-added chain. Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital. The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted [^7] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961).. By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
$Y=\Delta K+\Delta H+C_{p}\ ,\quad\text{allowing human capital,}$, (B.2)
@@
$C=C_{s}+C_{p}\quad\text{and}\quad\Delta H=C_{s}+W_{s}-D(H)\ ,\quad\text{allowing human capital.}$. (B.3)
@@
$Y=\Delta K+C+W_{s}-D(H)\ ,\quad\text{allowing human capital,}$, (B.4)
@@ Appendix C.  Mill’s statement of the free growth idea
This passage may be the first clear statement of what we call free growth theory. His use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.

### v11

@@ 1 Introduction and overview
Many economists over the centuries have reasoned that net saving, or equivalently net investment [^1] As reported in national accounts; they differ only by statistical discrepancy. Also see Appendix A., should tend to give equal capital growth. Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)). Roy Harrod (1939)) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”. Harrod, and many other economists of his time and since, have focused on growth of output rather than of capital, but have modeled growth of output by first assuming the equivalence of net saving and capital growth, within the warranted rate, and then looking for effects of that capital growth on later output growth.
@@
Those test results from data for net saving/capital ratio are easiest to reconcile in an interpretation that net saving has no effect on capital growth. [^2] Assuming that gross saving is enough to offset depreciation (see Section 3). We will assume this coverage unless stated otherwise. Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation. Value is created in the mind of the market at the moment when prospective cash flows are discounted. It is created only if the market sees a path, step by step, from the start, to practical realization of those prospective cash flows. Then capital growth arrives when the market first evaluates prospective cash flows, and is realized eventually in physical outcomes insofar as the market has predicted correctly. That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment. [^3] Present value is time-discounted, and affords innovators only enough, at first, for first steps in implementation. As implementation proceeds, present value rises and affords more. Early investors borrow from the future, in effect, and pay the interest from increments in present value as positive cash flows approach with time.
Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up. It is that temporary market advantage to the innovator which explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the source of investment. This idea will be called “free growth theory” for easy reference.
It predicts only at the largest scales, and only for the private sector. Individuals and groups and even small economies can grow in capital through saving if others dissave equally. That possibility is foreclosed only at the scale of all capital and all economies together. The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
@@
Inferences for economic teaching include the obvious ones for growth theory and for net saving in general. They include others as well. One of the central doctrines of the marginalist revolution has held that market realization converges to producer cost, when that cost includes imputed interest on assets owned. Net saving gives producer cost, and falls short of market realization in the presence of technological growth from new ideas. Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but not the aggregate value.
@@ 2 Net saving, capital growth and capital acceleration
$\Delta g(K)=\Delta s^{*}\ ,\quad\text{uta.}$ (3)
$\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative or zero. Division of Eq. (3) by capital acceleration, with rearrangement, gives
@@
Test results for the predictions $\frac{s^{*}}{g(K)}=\theta=1$, under thrift assumptions, appear in Fig. 1 and in Tables 1, 2 and 3. Screens to which they refer are explained in Section 14 below. Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts. Figures 1 and 2 show $\frac{s^{*}}{g(K)}$ and $\frac{\Delta s^{*}}{\Delta g(K)}$ as 0.481 and 0.064 respectively. Table 1 shows regression of $s^{*}$ on $g(K)$ as 0.0771, and regression of $\Delta s^{*}$ on $\Delta g(K)$ as 0.0559.
@@
Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗
Observations 1,826
R2 0.84511
Within R2 0.10669
Year fixed effects $\checkmark$
Country fixed effects $\checkmark$
Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗
Observations 1,574
R2 0.37071
Within R2 0.27233
Year fixed effects $\checkmark$
Country fixed effects $\checkmark$
@@
Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$. See Section 3 for predictions of free growth theory.
@@ 3 Interpretation of test results
Free growth theory predicts that net saving contributes nothing to capital growth, given that gross investment offsets depreciation. In years of recession, households may defer part or all of depreciation saving to support consumption. In these years, growth in the present value of better ideas will not realize equal net growth of capital, as some of it must offset any shortfall in depreciation saving first. Allowing for these years, we can say more compactly that free growth theory predicts that a gain in the present value of new ideas accounts for all capital creation, and thus for the sum of capital growth and restoration of shortfalls in depreciation investment. Thus free growth theory does not predict a zero value for $\theta$, but rather predicts $\theta$ to equal the share of capital creation applied to restoring those shortfalls. Insofar as 0.064 gives this share, free growth theory could fit the appearance in Fig. 2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises. Pending clarification of that share, we will proceed on the simplifying assumption that free growth theory predicts correctly.
@@ 5 Solow’s puzzle
Some growth is capital widening, where structures and implements increase in number but do not change in design. Capital widening, however, is practical only so far before glut and diminishing returns set in. Further growth must come from capital deepening, meaning improvements in the design of capital. Solow (1956) noted a kind of middle ground between capital widening and capital deepening in the disembodied growth mentioned earlier; ships carrying coals to Newcastle can raise prospective cash flows, and hence present value, by reversing the business plan. But Solow, who came to conclusions similar to ours from different evidence, puzzled as to how capital growth without net saving could be possible for capital deepening through “embodied” growth, where products of new design are made from plant of new design.
Solow received the Nobel Prize in 1987. His acceptance speech mentions the embodiment problem: [^5] Solow (1988). The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
@@ 6 Mechanics of free growth
The present value argument just given offers an accounting explanation of embodied growth in terms of sources of funds. The step-by-step mechanical explanation, we suggest, in terms of things rather than dollars, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.666i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve, and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
Our findings support those of Piketty and Zucman (2014) and Kurz (2023) as to the market power of innovators to explain capital growth beyond net saving. Those studies assumed that net saving is recovered in capital growth, so that this market power, which we call the innovator’s reserve, need explain only the part of capital growth remaining. Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
@@ 11 Net output
The concept of net output is found in our discussions, but not in our equations. Net output means value added. Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital. Both of these flows are unmeasurable by known means. Since we work here from measurable data for net saving and market value capital, we did our best to reason from those two quantities alone.
@@ 15 Disclaimers
Evidence supports the proposal of Modigliani (1954) that the young save to support retirement, and that retirees dissave to support consumption. [^8] Friedman (1957) follows a different argument to a similar conclusion. Such age-based differences are expected to offset in the stationary state, where tradition and free growth theory agree that overall net saving is zero. We accept that individual net saving can realize equal individual capital growth, so long as collective saving and dissaving by all individuals offset to zero.
@@ Appendix B. Net output with human capital
Human capital is impractical to measure, as it leaves little market record other than for its rental income in pay and investment cost in nurture and schooling. Thus national accounts leave it implicit, and allow us to infer what we can from data for pay and nurture and schooling. Those accounts are founded on the principle, sound when terms are appropriately redefined, that net output, or value added, is expressed in the sum of capital growth and net outflow from the value-added chain. In national accounts, then, where physical capital is the whole of capital while net outflow of the chain is the whole of consumption, the reasoning is
$Y=\Delta K+C\ ,\quad\text{neglecting human capital.}$ (B.1)
It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq. (B.1). Let human capital $H$, in that new model, stand as the last link in the value-added chain. Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital. The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted. [^9] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961). By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
$Y=\Delta K+\Delta H+C_{p}\ ,\quad\text{allowing human capital,}$ (B.2)
@@
$C=C_{s}+C_{p}\quad\text{and}\quad\Delta H=C_{s}+W_{s}-D(H)\ ,\quad\text{allowing human capital.}$ (B.3)
@@
$Y=\Delta K+C+W_{s}-D(H)\ ,\quad\text{allowing human capital,}$ (B.4)
@@ Appendix C. Mill’s statement of the free growth idea
This passage may be the first clear statement of what we call free growth theory. Mill’s use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.